Sol Summary – March 29th

Crew 177 Sol 4 Summary 29Mar2017
Person filling out Report:
Pitchayapa Jingjit—a biologist and a photographer for Crew 177
Summary Title:
Sol 4 at Mars
Mission Status:
Everyone is doing great and is hopeful about the toilet situation
SOL Activity Summary:
Improving individual/group research project and continuing the chemistry EVA. Depending on tonight’s sky, astronomy EVA may be conducted.
Look ahead Plan:
Continuing the geology EVA at the Little Canyon and depending on tonight’s sky, the astronomy EVA may be performed, and progressing through with individual/group projects.
Anomalies in Progress:
The weather is sunny in the morning and afternoon. We are expecting a clear sky tonight.
Crew Physical Status:
Everyone is better adjusted at Mars and is happy that the plumber came to the HAB.


EVA Report – March 29th

EVA Report:  EVA #6
Crew members: Elijah Espinoza, Caleb Li, Esteban Espinoza, Becky Parker
Location: 12 N 520446 4251115 (Candor Chasma)
Time: departure at 9:00 am. Estimated return at 11:30 am
Duration: 2 hours and 30 minutes approximately
Vehicles: 2 ATV’s and 1 Rover
Chemistry EVA.  Students will be visiting the location known as Candor Chasma with the goal of collecting samples as well collect water sample.  All samples will be analyzed in the Science Lab paying close attention to the chemical composition and Teach students about using the Geiger counter to identify background radiation.  Also have students collect a water sample for analysis back at the hab.
Goal: Complete the Chemistry EVA.
Unable to use the Geiger counter due undetermined technical issues. No water was found. For fauna, we observed birds, and found what appeared to be several nests, probably by a small mammal. Flora observed were flowers, grass, trees, cacti and brush. There were also many varieties of gypsum.  Some rock samples were brought back to the Hab for future chemical composition analysis.
Otsmar Villarroel
Crew 177 Commander

Commander Report – March 29th

Commander Report
Sol 5
March 29,2017
Otsmar Villarroel, Crew 177 Commander
Although today’s temperature was slightly higher than previous days, the day here on Mars has been beautiful with a spectacular blue sky.
All six crewmembers were up by 0700 local time as requested by the crew Commander the night before.  The crew enjoyed together scrambled eggs with sausage and biscuits for breakfast.  A few minutes after breakfast, a group consisting of Espinoza, Espinoza, Li, and Parker was sent out to Candor Chasma for a second attempt of looking for the presence of water.
 While during this effort, as the first one few days ago, the presence of water was not detected, the crew found the presence of moisture in some areas by Candor Chasma.  It was noticed after digging into the sand in the bottom of the chasma. Also, the presence of moisture was observed underneath some rocks.  Results are promising and further research regarding this topic has to be done.
After 2 hours 30 minutes, approximately, by noon (Martian time) the expediters returned to the Hab.  All eight crewmembers enjoyed lunch together.  Lunch was prepared crew members Quaas and Jingjit.
Part of the afternoon was used by the crew members to work on their independent projects.  Crew members have found some difficulties with their projects and they are making plans to solve them as soon as possible before rotation is over.  Also, crew member Jingjit continued working on her microbiology experiments.
As usual, dinner time was used to spend time together and discuss tomorrow’s plans.  Today it is crew member LaBarre’s birthday and we celebrated it with a birthday cake. The crew is planning to complete the geology EVA, pending approval. This would be the second half of EVA #4 that was shortened due to NPR reporter.

Journalist Report – March 29th

Crew 177 Journalist Report 29Mar2017
Victoria LaBarre, XO/Chemist
This morning the crew woke up a little later than expected because we all stayed up late looking at the stars. The view was spectacular, and the crew was thrilled to see different constellations so clearly, such as the little dipper, Ursa Major, Leo and Lynx. Tonight, our astronomers plan to re-calibrate the telescope so that we can see Jupiter.
Today we only sent out four crew members on a single EVA to Little Canyon to study geological features and collect dirt samples to test in the lab later.  Measuring the canyon proved difficult with the measuring tape since it was too short, so our geologist plans to return to the canyon on a later EVA with a rope that he can mark out on EVA and measure back at the Hab.
After lunch, our crew regrouped and assessed our projects and how to fix the ones that were not working. After this meeting, crew members went downstairs and worked on their projects, reaping positive results. Our microbiologist went around the lab swabbing random objects to grow bacteria on an auger, our greenhouse keeper got his bike generator working again and was able to run two tests, and the two crew members working on the robot soldered together a USB to USB cord made from two cut up charger cords and were able to send gain a connection between the robot’s brain and controller. The robot now can move its axels and arm, although it cannot move the heavy wheels.
The two crew members hypothesize that the 9 volt batteries do not feed enough power to the gearboxes to move the gears, chain, sprockets and wheel altogether. On a positive note, the arm of the robot is fully functional and these two crewmates plan to go on an EVA later to test the strength and dexterity of the robot’s arm.
Dinner was also very eventful as it was the OX/Journalist’s twenty-first birthday. The crew made a birthday cake with candles on top and all signed a birthday card with well wishes. Later, the crew broke out the glow sticks and looked at the stars again.
Unfortunately, the astronomers could not get the telescope to focus, so the distinguishing marks of Jupiter couldn’t be seen. However, glow sticks, a cup of hot cocoa and lots of laughs as we all looks up at the stars more than made up for the telescope.

Chemistry EVA – March 29th

Crew 177 Chemistry EVA 29Mar2017
Today, we analyzed the samples taken from Candor Canyon. After taking two 2 gram samples from each main rock sample, we used Hydrochloric Acid (HCl) to test for the presence of carbonate in the rock samples.
Separately, we used Nitric Acid (HNO3) and Silver Nitrate (AgNO3) in order to test for the presence of halides in the rock samples. The results showed positive results by showing that both carbonates and halides were present in the rock samples.
The positive carbonate (CO3-2) results are significant because it shows the presence of a potential atmosphere-building gas in Martian rock. In addition, the halide tests were significant because halides can be refined and used for a thermal-based fuel. The presence of halides in Martian rock shows that there is a potential energy source on Mars should solar and nuclear power methods fail.
Thank you,
Joseph Quaas (crew chief, crew chemist)

Geologist & HSO Report – March 28th

Crew 177 Geologist & HSO Officer


This is the third day being this HAB. I still cannot get rid of the excited feeling. Today is also the media day after we woke up in the morning around 7 am, we had the interview with the National Public Radio journalist, Ms. Rae Bichell. In the morning, four people from our crew and the reporter went out for a short EVA at the Cow Patty Field. In the afternoon, Dr. Villarroel, Joseph Quass, Victoria LaBarre and I went out to the Geology EVA. We ride 2 ATVs and 1 Rover, we went to the two stops, first stop is Cow Patty field, and we collected white-coating rock sample from the geology site 1A and 1B, sketched the sedimentary structure and took some photos of the field. We spent 50 mins on this stop. The next stop is the Little Canyon, we measured the height of the canyon, collected 8 bags of the dirt samples at the canyon, and took some photos there. We spent around 50 mins. Then, we came back to the HAB to enjoy the delicious tortilla meal.


Crew health situation: Everyone feels good, but we are worried about the lack of toilet and it might affect health.


Caleb Li

EVA Report – March 28th

EVA # 5

Filled by Otsmar Villarroel, Commander

Crew members:  Otsmar Villarroel, Caleb Li, Victoria LaBarre, Joseph Quass


12 N 518524 4249129
12 N 519235 4251489

Time: departure at 2:30 pm. Estimated return at 4:30 pm

Duration: 2 hours  approximately
Vehicles: 2 ATV’s and 1 Rover


Geology EVA.  Teach students about sedimentary rocks, deposition, lithification, and examination of the regolith.  Perform Monsell Soil Color analysis, use Brunton Compass to determine strike and dip.

Goal: Complete the Geology EVA.

EVA departure was slightly delayed due to the presence of the NPR reporter, Rae Bichell. During the EVA, we examined pieces of sandstone and used a Brunton compass to obtain the strike and dip of the cap rock on a hoodoo located at the site. Everyone was able to learn and demonstrate how to use the compass. The whole EVA was successfully completed.

EVA Report – March 28th

EVA # 4

Filled by Otsmar Villarroel, Commander

Crew members:  Elijah Espinoza, Esteban Ramirez, Pitchayapa Jingjit, Becky Parker

12 N 518524 4249129
12 N 519235 4251489

Time: departure at 10:25 am. Estimated return at 12:00 am

Duration: 1 hours and 35 minutes approximately
Vehicles: 2 Rovers, 1 ATV

Geology EVA.  Teach students about sedimentary rocks, deposition, lithification, and examination of the regolith.  Perform Monsell Soil Color analysis, use Brunton Compass to determine strike and dip.

Goal: Complete the Geology EVA.

EVA departure was slightly delayed due to the arrival of the NPR reporter, Rae Bichell. During the EVA, we examined pieces of sandstone and used a Brunton compass to obtain the strike and dip of the cap rock on a hoodoo located at the site. Everyone was able to learn and demonstrate how to use the compass. Rea accompanied the crew during the EVA, so we decided to made just one stop (12 N 518524 4249129).  Second part of the EVA will be rescheduled to be completed later this week.

Commander Report – March 28th

Four sols have passed since The McLennan Community College Highlanders Crew 177 arrived to Mars.  For sure, four sols full of challenges and amazing experiences.  Every activity around here is designed with the main purpose of providing us a unique experience other than that offered by a traditional classroom environment.

After all eight crewmembers took breakfast together, at about 0845 local time; Crew 177 got visited by Rae Bichell from NPR.  Rae Bichell spent most of the day with us learning about our routine here on Mars. She is was very interested in knowing about our experiences, projects, our college as well as all related with the MDRS program.  After giving her a short talk about our goals followed by a tour around the MDRS complex, she was invited to escort us during our morning Geology EVA.

After the daily routine of getting suited followed by the depressurization  process before going out of the Hab, the Geology EVA crew departed at about 1045 local time.  The EVA group consisted of Elijah Espinoza, Pitchayapa Jingjit, Esteban Espinoza, and Becky Parker. The geology experiments conducted today consisted of assessing many different regolith characteristics in the area around the Hab. Understanding what this material is made out of helps to determine what resources could be available. Also, crew members learned about the use of a Brunton compass to measure as strike and dip, both useful geological parameters which provide valuable information about what angle a particular geological formation has been uplifted.

While EVA team were collecting samples and taking geological measurement outside the Hab, remaining crew stayed in the Hab serving as communication support.  During that time they also prepared lunch which was enjoyed by the crew members and guest.  Menu included delicious teriyaki chicken, rice, corn, and red pepper.

During the afternoon, Pitchayapa started working on her microbiology experiments.  The point of her experiments is to swab the Hab for the presence of Gram negative and enteric bacteria and Gram positive staph bacteria.  Both have health implications.

Engineer completed their daily checking routine without major anomalies.

During dinner time the crew discussed the plan for tomorrow.  While some crew members showed some signs of fatigue during the day, Crew 177 members are looking forward to keep doing their best in order to successfully complete this mission.  The challenges make us stronger!
Tomorrow, pending approval, will consist of a morning chemistry EVA with Elijah, Esteban, Caleb, and Becky.  Rest of the afternoon will be spend working on individual projects and microbiology experiments conducted by Pitchayapa.

Good night from Mars